eClipSe:

If x,y,z are real numbers such that #x^2+y^2+2z^2=4x-2z+2yz-5# then #x-y-z=?#

xXMarcelieXx:

what math is this?

nuts:

\(x^2+y^2+2z^2=4x-2z+2yz-5\) \(x^2-4x+4+y^2+2z^2=-2z+2yz-1\) \((x+2)^2+y^2-2yz+z^2+z^2=-2z-1\) \((x+2)^2+(y-z)^2+z^2=-2z-1\) \((x+2)^2+(y-z)^2+z^2+2z+1=0\) \((x+2)^2+(y-z)^2+(z+1)^2=0\) \(x=-2\) \(y-z=0\) \(z=-1\) \(x=-2, y=-1, z=-1\) \(x-y-z=-2+1+1=0\)